cullenn:

Hey can someone help me with this ixl? M.5 Similar triangles and indirect measurement

cullenn:

1 attachment
DUXK:

15 meters is the answer

cullenn:

@duxk wrote:
15 meters is the answer

1 attachment
DUXK:

q would be 76

iluvpanda:

You should teach him how to do it not just give him the answers

DUXK:

@iluvpanda wrote:
You should teach him how to do it not just give him the answers
alr

DUXK:

The two triangles POQ and QMN are both similar and the marked angles show they are equal or AA similarity. PO = 44ft, MN = 33ft, PN = 57ft and PQ = q 44/33 = q/57 and simplify it too 44/33 = 4/3 and then 4/3 = q/57 and then multiply both sides by 57. q = 57 x 4/3 and u get q = 19 x 4 and finally u get q = 76

cullenn:

@iluvpanda wrote:
You should teach him how to do it not just give him the answers
her-

iluvpanda:

OMG MB 😭😭

cullenn:

@duxk wrote:
The two triangles POQ and QMN are both similar and the marked angles show they are equal or AA similarity. PO = 44ft, MN = 33ft, PN = 57ft and PQ = q 44/33 = q/57 and simplify it too 44/33 = 4/3 and then 4/3 = q/57 and then multiply both sides by 57. q = 57 x 4/3 and u get q = 19 x 4 and finally u get q = 76

1 attachment
cullenn:

sob i think i understand it a little better js a few more exmaples-